# Practical and efficient quantum circuit synthesis and transpiling with Reinforcement Learning

### v1

@@ I Introduction
In terms of AI applied to quantum circuits, the literature offers several examples in the areas of synthesis [27, 28], optimization [29, 30, 31], mapping [32, 33] or compilation [34, 35, 36, 37, 38]. This paper builds upon existing state-of-the-art procedures and focuses on applying AI techniques to quantum circuit transpilation and optimization.
@@ III Discussion
There exist examples in literature of various approaches to applying artificial intelligence to quantum circuits. However, the state-of-the-art in this field is still limited. In many cases, existing work relies on existing datasets to train the algorithms [27, 28, 32, 45, 33, 37]. Taking as example [16], we observe the creation of a dataset with enough quality and scale (in terms of N of qubits, gates used, etc.) can be extremely large and expensive to compile or generate. This could be a barrier for many problems, as without massive datasets, some AI-based approaches will not be useful or will not scale to solve quantum-circuit-related tasks involving hundreds or even tens of qubits. In contrast, other authors are already exploring the use of other AI-based techniques (RL, genetic algorithms, etc.) where the existence of such datasets is not required [29, 30, 31, 34, 35, 36, 38]. This differentiation is relevant since although standard machine learning algorithms cannot surpass the data-generation process, RL (and similar approaches) demonstrate their strength by learning from experience, and bypass the need for annotated datasets. Continuing with the existing literature, we also observe a lack of device awareness in most of the existing works, meaning they enhance or solve different tasks involving quantum circuits but still require further transpiling to adapt to native instruction sets and connectivity restrictions of most current quantum devices.
@@
The combination of quality and speed make this method a strong candidate to be used in a real quantum computing environment. The approach described here is the foundation on which the Qiskit Transpiler Service [^2] https://docs.quantum.ibm.com/transpile/qiskit-transpiler-service, and its AI transpiler passes [^3] https://docs.quantum.ibm.com/transpile/ai-transpiler-passes are built. The service and its AI features are currently used by IBM Quantum users in real quantum computing workflows with positive results and a significant reduction of CNOTs and layers of CNOTs for most of the circuits.
We have shown different results for each type of circuit. For Linear Function circuits we present results up to 9 qubits, for Clifford up to 11 qubits, for Permutation up to 65 qubits and we discuss how we run routing up to 133 qubits. Each type of circuit has their own actions space and particularities, and our approach scales differently for them. We acknowledge that the synthesis methods, as presented here, could struggle to scale to thousands of qubits, as we have seen that the training procedure takes longer to converge when increasing the circuit size. As far as we know, there is no fundamental limitation on scaling the method -the size of the networks might in principle scale linearly with the size of the operators and the gate sets-, so improving the method to train more efficiently on larger circuits is an interesting research direction. However, we argue that, in practice, the method is useful even with the circuit sizes presented here, since we use this as an optimization stage for re-synthesizing certain blocks within larger circuits.
@@ Acknowledgements.
We would like to express our gratitude to the Data and Intelligence team at IBM Quantum for their efforts in implementing the procedures presented in this paper within the Qiskit Transpiler Service. We specifically acknowledge Jesús Talavera, Sanjay Vishwakarma, and Yaiza García Martín Mantero for their invaluable contributions to this project. The authors acknowledge feedback and insightful discussions with Ali Javadi, Blake R. Johnson, Lev Bishop, Jay Gambetta, and Francisco J. Martín-Fernández. The authors also acknowledge the IBM Research Cognitive Computing Cluster service for providing resources that have contributed to the research results reported within this paper.

### v2

@@ I Introduction
In terms of AI applied to quantum circuits, the literature offers several examples in the areas of synthesis [27, 28, 29], optimization [30, 31, 32, 33], mapping [34, 35] or compilation [36, 37, 38, 39, 40, 41]. This paper builds upon existing state-of-the-art procedures and focuses on applying AI techniques to quantum circuit transpilation and optimization.
@@ III Discussion
There exist examples in literature of various approaches to applying artificial intelligence to quantum circuits. However, the state-of-the-art in this field is still limited. In many cases, existing work relies on existing datasets to train the algorithms [27, 28, 34, 48, 35, 39]. Taking as example [16], we observe the creation of a dataset with enough quality and scale (in terms of N of qubits, gates used, etc.) can be extremely large and expensive to compile or generate. This could be a barrier for many problems, as without massive datasets, some AI-based approaches will not be useful or will not scale to solve quantum-circuit-related tasks involving hundreds or even tens of qubits. In contrast, other authors are already exploring the use of other AI-based techniques (RL, genetic algorithms, etc.) where the existence of such datasets is not required [30, 31, 32, 36, 37, 38, 40, 29, 33, 41]. This differentiation is relevant since although standard machine learning algorithms cannot surpass the data-generation process, RL (and similar approaches) demonstrate their strength by learning from experience, and bypass the need for annotated datasets. Continuing with the existing literature, we also observe a lack of device awareness in most of the existing works, meaning they enhance or solve different tasks involving quantum circuits but still require further transpiling to adapt to native instruction sets and connectivity restrictions of most current quantum devices.
@@
The combination of quality and speed make this method a strong candidate to be used in a real quantum computing environment. The approach described here is the foundation on which the Qiskit Transpiler Service [^2] https://docs.quantum.ibm.com/guides/qiskit-transpiler-service, and its AI-powered transpiler passes [^3] https://docs.quantum.ibm.com/guides/ai-transpiler-passes are built. The service and its AI features are currently used by IBM Quantum users in real quantum computing workflows with positive results and a significant reduction of CNOTs and layers of CNOTs for most of the circuits.
We have shown different results for each type of circuit. For Linear Function circuits we present results up to 9 qubits, for Clifford up to 11 qubits, for Permutation up to 65 qubits and we discuss how we run routing up to 133 qubits. Each type of circuit has their own actions space and particularities, and our approach scales differently for them. We acknowledge that the synthesis methods, as presented here, could struggle to scale to thousands of qubits, as we have seen that the training procedure takes longer to converge when increasing the circuit size. As far as we know, there is no fundamental limitation on scaling the method -the size of the networks might in principle scale linearly with the size of the operators and the gate sets-, so improving the method to train more efficiently on larger circuits is an interesting research direction. However, we argue that, in practice, the method is useful even with the circuit sizes presented here, since we use this as an optimization stage for re-synthesizing certain blocks within larger circuits. These statements presented about the potential of our solution with RL in terms of scalability, applicability in real contexts, and competitiveness in the balance of efficiency and results have been independently verified in studies such as [49] where the authors have been able to execute benchmarks that show how our AI methods included in the Qiskit Transpiler Service are able to scale to circuits of up to 900 qubits, and one million two-qubit gates, yielding results that are at least $\sim$30better in terms of depth and two-qubit gates count on average compared to other quantum computing libraries and their heuristic algorithms.
@@ Acknowledgements.
We would like to express our gratitude to the Quantum+AI team at IBM Quantum for their efforts in implementing the procedures presented in this paper within the Qiskit Transpiler Service. We specifically acknowledge Jesús Talavera, Sanjay Vishwakarma, and Yaiza García Martín Mantero for their invaluable contributions to this project. The authors acknowledge feedback and insightful discussions with Ali Javadi, Blake R. Johnson, Lev Bishop, Jay Gambetta, and Francisco J. Martín-Fernández. The authors also acknowledge the IBM Research Cognitive Computing Cluster service for providing resources that have contributed to the research results reported within this paper.